conducted to determine success rates of students enrolled in the Statistics courses offered at South Plains College for the fall semester of 2015. A random sample of 14 students was taken, and we recorded each student's age, final average, gender, # hours worked per week, race, and t attended this semester). Use the results from the following table to answer all parts of #2. he attendance record (# of classes not Age 20 18 19 Final Average Gender # Hrs worked/wk 25 35 40 25 40 Race Attendance 82 50 68 95 6 19 43 24 25 50 18 17 64 89 86 76 81 20 50 32 15 0 20 21 26 20 70 78 30 40 20 84
If three students are randomly selected, find the probability that at least two is a male. (3 pts) k. Find the 99% confidence interval for the mean age of all Statistics students this semester. (3 pts) I.
If3 students are randomly selected, what is the probability that all will be females? (3 pts) m. n. We want to estimate the percentage of students next semester who earned a final average of 90 or above. How many students mustwe survey next semester, if we want to be 95% confident that our proportion estimate will be within 3 percentage points? (4 pts)
Answer #1

SOLUTION

1. In the above sample of 14 students, there are 5 males and 9 females. So, probability of selecting atleast one male = 1 - probability of selecting 0 males = 1 - 9C3/14C3 = 0.7692

2. 90% confidence interval for mean age: (population SD is unknown)

Sample mean +- t0.05,n-1* sample sd/√n

= 24.286 +- 1.96*9.895/√14

= (19.102, 29.469)

3. If 4 students are randomly selected, probability of all females = 9C3/14C3= 0.2307